if a car passes four station with a constant acceleration the station are named a,b,c,and d initially at rest if Vi=2Vb, Vd=3Vc and the distance traveled from B to c is 200m. find the total distance traveled?  


1
Expert's answer
2021-01-22T15:05:03-0500

The condition of the problem is incorrect.



We will accept that "v_a=0", "v_c=2v_b", "v_d=3v_b".


"a=const" and "BC=200 (m)".


"s_2=\frac{(2v_b)^2-v_b^2}{2a}\to a=\frac{3v_b^2}{2s_2}";


"s=\frac{v_d^2-v_a^2}{2a}=\frac{v_d^2-0}{2\cdot3v_b^2}\cdot2s_2=\frac{(3v_b)^2-0}{2\cdot3v_b^2}\cdot2\cdot200=600(m)". Answer.